5- Show work, answer the following for the given quadratic function.
F(x)=x²-4x+4
a- What is the vertex (h,k) of f?
Simplify your answer(type an ordered pair)
b- What is the axis of symmetry?
The axis of symmetry is
Type an equation simplify answer, use integers or fractions for any numbers in the equation
c- What are the interepts?
The intercepts are?
Type an ordered pair, use a comma to separate answers as needed
d- What is the domain of f?
e- What is the range of f?
f- On what interval is f increasing?
g- On what interval is f degreasing?
h- Graph

Answer by ewatrrr(24785)

 
Hi,
vertex(2,-8) and line of symmetry x = 2 and intercepts shown
Range: F(x) ≥ -8, Domain: x all real numbers
f Increaseing ( 2,∞ ) and f is decreasing(−∞,2)
F(x) = x^2 - 4x- 4 = x^2 - 4x + 4 - 4 - 4 = (x-2)^2 - 8